# Improved Regret Bounds for Gaussian Process Upper Confidence Bound in Bayesian Optimization

### v2

@@ B.2 Summary of Mercer Decomposition for Dot-Product Kernel on Sphere
In this subsection, we summarize the basic known results of the Mercer decomposition on $\mathbb{S}^{d}$. The content of this subsection is related to the analysis of the spherical harmonics. We refer to [1, 15] as the basic textbook. In the kernel method literature, the Mercer decomposition of the dot-product kernel and its eigendecay have been studied. See, e.g., [2, 33, 45]. Furthermore, the existing analysis of the neural tangent kernel also leverages the Mercer decomposition based on the spherical harmonics. We also refer to the appendix of [3, 5] as the related works of this subsection.
We first describe Mercer’s theorem. Let $L^{2}(\mathcal{X},\mu)\coloneqq\{f:\mathcal{X}\rightarrow\mathbb{R}\mid\int_{\mathcal{X}}f^{2}(\bm{x})\mu(\text{d}\bm{x})\}$ be the square-integrable functions on $\mathcal{X}$ under the measure $\mu$. Furthermore, let us define the kernel integral operator $\mathcal{T}_{k}:L^{2}(\mathcal{X},\mu)\rightarrow L^{2}(\mathcal{X},\mu)$ of a square-integrable kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ as $(\mathcal{T}_{k}f)(\cdot)=\int_{\mathcal{X}}k(\cdot,\bm{x})f(\bm{x})\mu(\text{d}\bm{x})$. Then, Mercer’s theorem guarantees that the positive kernel $k$ is decomposed based on the eigenvalues and eigenfunctions sequence of $\mathcal{T}_{k}$ with absolute and uniform convergence on $\mathcal{X}\times\mathcal{X}$. We give the formal statement below.
Theorem 10 (Mercer’s theorem, e.g., Theorem 4.49 in [12]).
@@
Definition 1 (Spherical harmonics, e.g., Definition 2.7 in [1]).
Fix any $d\geq 1$ and $m\in\mathbb{N}$. Let $\mathbb{Y}_{m}(\mathbb{R}^{d+1})$ be the all homogeneous polynomials of degree $m$ in $\mathbb{R}^{d+1}$ that are also harmonic [^11] A polynomial $H(x_{1},\ldots,x_{d+1})$ is called homogeneous of degree $m$ if $H(tx_{1},\ldots,tx_{d+1})=t^{n}H(x_{1},\ldots,x_{d+1})$. Furthermore, a polynomial $H(x_{1},\ldots,x_{d+1})$ is called harmonic if $\Delta_{d+1}H=0$, where $\Delta_{d+1}$ is the Laplace operator. See Chapter 4 in [15] or Chapter 2 in [1].. The space $\mathbb{Y}_{m}^{d+1}=\mathbb{Y}_{m}(\mathbb{R}^{d+1})\mid_{\mathbb{S}^{d}}$ is called the spherical harmonic space of order $m$ in $d+1$ dimensions. Any function in $\mathbb{Y}_{m}^{d+1}$ is called a spherical harmonic of order $m$ in $d+1$ dimensions.
@@
Lemma 11 (Dimension and completeness of sphererical harmonics, e.g., Chapter 2.1.3, Corollary 2.15, and Theorem 2.38 in [1]).
@@
Lemma 12 (Funk-Hecke Formula, e.g., Theorem 2.22 in [1] or Theorem 4.24 in [15]).
@@
Lemma 12 suggests that the spherical harmonics are eigenfunctions of the continuous dot-product kernel $k(\bm{x},\widetilde{\bm{x}})=\widetilde{k}(\bm{x}^{\top}\widetilde{\bm{x}})$ on $\mathbb{S}^{d}$. Furthermore, Lemma 11 guarantees the $\cup_{m\in\mathbb{N}}(Y_{m,j})_{j\in[N_{d+1,m}]}$ forms an orthonormal bases of $L^{2}(\mathbb{S}^{d},\sigma)$, which implies that they are the orthonormal bases of the eigenspace of $\mathcal{T}_{k}$. These facts give the following explicit form of Mercer decomposition for a continuous dot-product kernel.
@@
Fix any $d\in\mathbb{N}_{+}$. Suppose $\mathcal{X}=\mathbb{S}^{d}$. Furthermore, assume the kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ is the positive definite kernel such that $\forall\bm{x},\widetilde{\bm{x}}\in\mathcal{X},k(\bm{x},\widetilde{\bm{x}})=\widetilde{k}(\bm{x}^{\top}\widetilde{\bm{x}})$ with some continuous function $\widetilde{k}:[-1,1]\rightarrow\mathbb{R}$. Then, we have the following Mercer decomposition of $k$:
$k(\bm{x},\widetilde{\bm{x}})=\sum_{m=0}^{\infty}\lambda_{m}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x})Y_{m,j}(\widetilde{\bm{x}}),$ (188)
@@
Lemma 14 (Addition theorem, e.g., Theorem 2.9 in [1] or Theorem 4.11 in [15]).
@@
$\forall\bm{x},\widetilde{\bm{x}}\in\mathbb{S}^{d},~{}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x})Y_{m,j}(\widetilde{\bm{x}})=\frac{N_{d+1,m}}{|\mathbb{S}^{d+1}|}P_{m,d}(\bm{x}^{\top}\widetilde{\bm{x}}),$ (190)
@@ B.3 Upper Bound of MIG with Mercer Decomposition
Lemma 15 (Adapted from [53]).
@@
The proof almost directly follows from [54], while a minor modification is required to deal with the unboundness of the eigenfunctions through the addition theorem. The same proof strategy is already provided in [29, 53] for analyzing the MIG of the neural tangent kernel on the sphere. Although our proof has no intrinsic change from their proof, we give the details below for completeness of our paper.
@@
We first decompose the kernel matrix as $\mathbf{K}(\mathbf{X}_{T},\mathbf{X}_{T})=\mathbf{K}_{\mathrm{head}}+\mathbf{K}_{\mathrm{tail}}$, where $[\mathbf{K}_{\mathrm{head}}]_{i,l}=\sum_{m=0}^{M}\lambda_{m}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x}_{i})Y_{m,j}(\bm{x}_{l})$ and $[\mathbf{K}_{\mathrm{tail}}]_{i,l}=\sum_{m=M+1}^{\infty}\lambda_{m}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x}_{i})Y_{m,j}(\bm{x}_{l})$. Then, as with the proof in [54], the MIG is decomposed as
@@
Based on the feature representation of the kernel, the first term is further bounded from above as follows (see [53, 54]):
@@
where the second inequality follows from $N_{M}\geq 1$. Regarding the second term, as with [54], we have
@@
where the first inequality follows from $\ln\det(A)\leq T\ln(\mathrm{Tr}(A)/T)$ for any positive definite matrix $A\in\mathbb{R}^{T\times T}$ (e.g., Lemma 1 in [54]). Then, from addition theorem (Theorem 14), we have
@@ B.4 Eigendecay of SE and Matérn Kernel
To obtain the explicit upper bound of Eq. (191), we need the upper bound of the eigenvalue in Eq. (189) under SE and Matérn kernel. Regarding SE kernel, several existing works have already studied it [33, 36]. We formally provide the following lemma from [33].
Lemma 17 (Eigendecay for $k=k_{\mathrm{SE}}$ on $\mathbb{S}^{d}$, Theorem 2 in [33]).
@@
Regarding Matérn kernel, we provide the upper bound of $\lambda_{m}$ for $\nu>1/2$ by extending the proof in [19], which studies $\lambda_{m}$ for Laplace kernel (Matérn with $\nu=1/2$). As with the proof in [19], we leverage the following lemma, which relates the spectral density of the kernel to $\lambda_{m}$.
Lemma 18 (Eigenvalues and spectral density, Theorem 4.1 in [35]).
Fix any $d\in\mathbb{N}_{+}$. Suppose that $k:\mathbb{R}^{d+1}\times\mathbb{R}^{d+1}\rightarrow\mathbb{R}$ is a positive definite, stationary, and isotropic kernel function on $\mathbb{R}^{d+1}$ such that $\forall\bm{x},\widetilde{\bm{x}},k(\bm{x},\widetilde{\bm{x}})=\Phi(\bm{x}-\widetilde{\bm{x}})$ for some function $\Phi(\cdot)$. Furthermore, suppose $\Phi(\cdot)$ is represented as
@@
for some function $\widehat{\Phi}$ such that $\forall a\geq 0,\widehat{\Phi}(a)\geq 0$ and $\int_{\mathbb{R}^{d+1}}\widehat{\Phi}(\|\bm{\eta}\|_{2})\mathrm{d}\bm{\eta}<\infty$. Then, there exists a function $\widetilde{k}:[-1,1]\rightarrow\mathbb{R}$ such that $\forall\bm{x},\widetilde{\bm{x}}\in\mathbb{S}^{d},\widetilde{k}(\bm{x}^{\top}\widetilde{\bm{x}})=k(\bm{x},\widetilde{\bm{x}})$. Furthermore, $\lambda_{m}$ in Eq. (189) is given by
@@
See, Chapter 4.2 in [38]. By using Lemma 18, we obtain the following lemma.
@@
As with the proof of Theorem 7 in [19], we evaluate the integral $\int_{0}^{\infty}t^{-2\nu-d}B_{m+\frac{d-1}{2}}^{2}(t)\mathrm{d}t$ by using the following identity (Chapter 13.4.1 in [20]):
@@ B.5 Proof of Theorem 7
where the first and last inequalities follow from Lemmas 9, 15 and Lemma 16, respectively. Here, for any $d\in\mathbb{N}_{+}$ and $m\in\mathbb{N}_{+}$, we have $m^{\frac{d}{2}-1}\leq c_{d}^{m}$ with $c_{d}=\max\left\{1,\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)\right\}$. Indeed, when $d\leq 2$, we have $m^{d/2-1}\leq 1=c_{d}$. When $d\geq 3$, the function $g(m)=m^{\frac{1}{m}\left(\frac{d}{2}-1\right)}$ attains maximum at $m=e$ on $[1,\infty)$, which implies $g(m)\leq\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)\Rightarrow m^{d/2-1}\leq\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)^{m}\leq c_{d}^{m}$. Hence, we have
@@ Appendix C Auxiliary Lemmas
Suppose $f$ is continuous. Then, the following two statements hold:
  • Under conditions 1 and 3 in Lemma 2, $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ holds for any $\bm{x}\in\mathcal{X}$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}$ with $\varepsilon_{1}=\min\{c_{\mathrm{gap}},c_{\mathrm{lin}}\rho_{\mathrm{lin}}\}$.
  • Under conditions 1 and 4 in Lemma 2, $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ holds for any $\bm{x}\in\mathcal{X}$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{2}$ with $\varepsilon_{2}=\min\{c_{\mathrm{gap}},c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}\}$.
@@
We only prove the first statement since the proof of the second statement is almost the same. When $\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})=\mathcal{X}$, the statement is trivial. Hereafter, we assume $\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})\neq\mathcal{X}$. Here , note that $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{lin}}\rho_{\mathrm{lin}}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ from condition 3 in Lemma 2, where $\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})=\{\bm{x}\in\mathcal{X}\mid\|\bm{x}-\bm{x}^{\ast}\|_{2}=\rho_{\mathrm{lin}}\}$. Furthermore, from the continuity of $f$ and the compactness of $(\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast}))\cup\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$, there exists $\widetilde{\bm{x}}_{\ast}\in\operatornamewithlimits{argmax}_{\bm{x}\in(\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast}))\cup\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})}f(\bm{x})$. Then, we consider the following two cases separately.
@@
  • When $c_{\mathrm{gap}}\geq c_{\mathrm{lin}}\rho_{\mathrm{lin}}$, $\varepsilon_{1}=c_{\mathrm{lin}}\rho_{\mathrm{lin}}$ holds. If there exists $\bm{x}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}=c_{\mathrm{lin}}\rho_{\mathrm{lin}}$, we can choose $\widetilde{\bm{x}}_{\ast}$ such that $\widetilde{\bm{x}}_{\ast}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ since $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{lin}}\rho_{\mathrm{lin}}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$. Furthermore, such $\widetilde{\bm{x}}_{\ast}$ is the local maximizer on $\mathcal{X}$, which satisfies $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}}_{\ast})\leq f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}\leq c_{\mathrm{gap}}.$ This contradicts condition 1 in Lemma 2.
@@
  • When $c_{\mathrm{gap}}<c_{\mathrm{lin}}\rho_{\mathrm{lin}}$, $\varepsilon_{1}=c_{\mathrm{gap}}$ holds. If there exists $\bm{x}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}=c_{\mathrm{gap}}$, we can choose $\widetilde{\bm{x}}_{\ast}$ such that $\widetilde{\bm{x}}_{\ast}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$ since $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{lin}}\rho_{\mathrm{lin}}>c_{\mathrm{gap}}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{lin}};\bm{x}^{\ast})$. Furthermore, such $\widetilde{\bm{x}}_{\ast}$ is the local maximizer on $\mathcal{X}$, which satisfies $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}}_{\ast})\leq f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}=c_{\mathrm{gap}}.$ This contradicts condition 1 in Lemma 2.
@@
[…]
@@
where the second inequality follows from Theorems 5.3 and 5.4 in [48]. ∎

### v3

@@ B.2 Summary of Mercer Decomposition for Dot-Product Kernel on Sphere
In this subsection, we summarize the basic known results of the Mercer decomposition on $\mathbb{S}^{d}$. The content of this subsection is related to the analysis of the spherical harmonics. We refer to [atkinson2012spherical, efthimiou2014spherical] as the basic textbook. In the kernel method literature, the Mercer decomposition of the dot-product kernel and its eigendecay have been studied. See, e.g., [azevedo2014sharp, minh2006mercer, scetbon2021spectral]. Furthermore, the existing analysis of the neural tangent kernel also leverages the Mercer decomposition based on the spherical harmonics. We also refer to the appendix of [bach2017breaking, bietti2020deep] as the related works of this subsection.
We first describe Mercer’s theorem. Let $L^{2}(\mathcal{X},\mu)\coloneqq\{f:\mathcal{X}\rightarrow\mathbb{R}\mid\int_{\mathcal{X}}f^{2}(\bm{x})\mu(\text{d}\bm{x})<\infty\}$ be the square-integrable functions on $\mathcal{X}$ under the measure $\mu$. Furthermore, let us define the kernel integral operator $\mathcal{T}_{k}:L^{2}(\mathcal{X},\mu)\rightarrow L^{2}(\mathcal{X},\mu)$ of a square-integrable kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ as $(\mathcal{T}_{k}f)(\cdot)=\int_{\mathcal{X}}k(\cdot,\bm{x})f(\bm{x})\mu(\text{d}\bm{x})$. Then, Mercer’s theorem guarantees that the positive kernel $k$ is decomposed based on the eigenvalues and eigenfunctions sequence of $\mathcal{T}_{k}$ with absolute and uniform convergence on $\mathcal{X}\times\mathcal{X}$. We give the formal statement below.
Theorem 10 (Mercer’s theorem, e.g., Theorem 4.49 in [christmann2008support]).
@@
Definition 1 (Spherical harmonics, e.g., Definition 2.7 in [atkinson2012spherical]).
Fix any $d\geq 1$ and $m\in\mathbb{N}$. Let $\mathbb{Y}_{m}(\mathbb{R}^{d+1})$ be the all homogeneous polynomials of degree $m$ in $\mathbb{R}^{d+1}$ that are also harmonic [^11] A polynomial $H(x_{1},\ldots,x_{d+1})$ is called homogeneous of degree $m$ if $H(tx_{1},\ldots,tx_{d+1})=t^{n}H(x_{1},\ldots,x_{d+1})$. Furthermore, a polynomial $H(x_{1},\ldots,x_{d+1})$ is called harmonic if $\Delta_{d+1}H=0$, where $\Delta_{d+1}$ is the Laplace operator. See Chapter 4 in [efthimiou2014spherical] or Chapter 2 in [atkinson2012spherical].. The space $\mathbb{Y}_{m}^{d+1}=\mathbb{Y}_{m}(\mathbb{R}^{d+1})\mid_{\mathbb{S}^{d}}$ is called the spherical harmonic space of order $m$ in $d+1$ dimensions. Any function in $\mathbb{Y}_{m}^{d+1}$ is called a spherical harmonic of order $m$ in $d+1$ dimensions.
@@
Lemma 11 (Dimension and completeness of sphererical harmonics, e.g., Chapter 2.1.3, Corollary 2.15, and Theorem 2.38 in [atkinson2012spherical]).
@@
Lemma 12 (Funk-Hecke Formula, e.g., Theorem 2.22 in [atkinson2012spherical] or Theorem 4.24 in [efthimiou2014spherical]).
@@
[…]
@@
[…]
@@
Lemma 14 (Addition theorem, e.g., Theorem 2.9 in [atkinson2012spherical] or Theorem 4.11 in [efthimiou2014spherical]).
@@
[…]
@@ B.3 Upper Bound of MIG with Mercer Decomposition
Lemma 15 (Adapted from [vakili2021uniform]).
@@
The proof almost directly follows from [vakili2021information], while a minor modification is required to deal with the unboundness of the eigenfunctions through the addition theorem. The same proof strategy is already provided in [kassraie2022neural, vakili2021uniform] for analyzing the MIG of the neural tangent kernel on the sphere. Although our proof has no intrinsic change from their proof, we give the details below for completeness of our paper.
@@
We first decompose the kernel matrix as $\mathbf{K}(\mathbf{X}_{T},\mathbf{X}_{T})=\mathbf{K}_{\mathrm{head}}+\mathbf{K}_{\mathrm{tail}}$, where $\mathbf{K}_{\mathrm{head}}]_{i,l}=\sum_{m=0}^{M}\lambda_{m}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x}_{i})Y_{m,j}(\bm{x}_{l})$ and $\mathbf{K}_{\mathrm{tail}}]_{i,l}=\sum_{m=M+1}^{\infty}\lambda_{m}\sum_{j=1}^{N_{d+1,m}}Y_{m,j}(\bm{x}_{i})Y_{m,j}(\bm{x}_{l})$. Then, as with the proof in [vakili2021information], the MIG is decomposed as
@@
Based on the feature representation of the kernel, the first term is further bounded from above as follows (see [vakili2021uniform, vakili2021information]):
@@
where the second inequality follows from $N_{M}\geq 1$. Regarding the second term, as with [vakili2021information], we have
@@
where the first inequality follows from $\ln\det(A)\leq T\ln(\mathrm{Tr}(A)/T)$ for any positive definite matrix $A\in\mathbb{R}^{T\times T}$ (e.g., Lemma 1 in [vakili2021information]). Then, from addition theorem (Theorem 14), we have
@@ B.4 Eigendecay of SE and Matérn Kernel
To obtain the explicit upper bound of Eq. (169), we need the upper bound of the eigenvalue in Eq. (167) under SE and Matérn kernel. Regarding SE kernel, several existing works have already studied it [minh2006mercer, narcowich2007approximation]. We formally provide the following lemma from [minh2006mercer].
Lemma 17 (Eigendecay for $k=k_{\mathrm{SE}}$ on $\mathbb{S}^{d}$, Theorem 2 in [minh2006mercer]).
@@
Regarding Matérn kernel, we provide the upper bound of $\lambda_{m}$ for $\nu>1/2$ by extending the proof in [geifman2020similarity], which studies $\lambda_{m}$ for Laplace kernel (Matérn with $\nu=1/2$). As with the proof in [geifman2020similarity], we leverage the following lemma, which relates the spectral density of the kernel to $\lambda_{m}$.
Lemma 18 (Eigenvalues and spectral density, Theorem 4.1 in [narcowich2002scattered]).
[…]
@@
[…]
@@
See, Chapter 4.2 in [Rasmussen2005-Gaussian]. By using Lemma 18, we obtain the following lemma.
@@
As with the proof of Theorem 7 in [geifman2020similarity], we evaluate the integral $\int_{0}^{\infty}t^{-2\nu-d}B_{m+\frac{d-1}{2}}^{2}(t)\mathrm{d}t$ by using the following identity (Chapter 13.4.1 in [gray1895treatise]):
@@ B.5 Proof of Theorem 7
where the first and last inequalities follow from Lemmas 9, 15 and Lemma 16, respectively. Here, for any $d\in\mathbb{N}_{+}$ and $m\in\mathbb{N}_{+}$, we have $m^{\frac{d}{2}-1}\leq c_{d}^{m}$ with $c_{d}=\max\left\{1,\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)\right\}$. Indeed, when $d\leq 2$, we have $m^{d/2-1}\leq 1=c_{d}$. When $d\geq 3$, the function $g(m)=m^{\frac{1}{m}\left(\frac{d}{2}-1\right)}$ attains maximum at $m=e$ on $1,\infty)$, which implies $g(m)\leq\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)\Rightarrow m^{d/2-1}\leq\exp\left(\frac{1}{e}\left(\frac{d}{2}-1\right)\right)^{m}\leq c_{d}^{m}$. Hence, we have
@@ Appendix C Auxiliary Lemmas
Suppose $f$ is continuous. Then, under conditions 1 and 3 in Lemma 2, $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ holds for any $\bm{x}\in\mathcal{X}$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon$ with $\varepsilon=\min\{c_{\mathrm{gap}},c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}\}$.
@@
When $\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})=\mathcal{X}$, the statement is trivial. Hereafter, we assume $\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})\neq\mathcal{X}$. Here, note that $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ from condition 3 in Lemma 2, where $\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})=\{\bm{x}\in\mathcal{X}\mid\|\bm{x}-\bm{x}^{\ast}\|_{2}=\rho_{\mathrm{quad}}\}$. Furthermore, from the continuity of $f$ and the compactness of $(\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast}))\cup\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$, there exists $\widetilde{\bm{x}}_{\ast}\in\operatornamewithlimits{argmax}_{\bm{x}\in(\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast}))\cup\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})}f(\bm{x})$. Then, we consider the following two cases separately.
@@
  • When $c_{\mathrm{gap}}\geq c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$, $\varepsilon=c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$ holds. If there exists $\bm{x}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon=c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$, we can choose $\widetilde{\bm{x}}_{\ast}$ such that $\widetilde{\bm{x}}_{\ast}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ since $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$. Furthermore, such $\widetilde{\bm{x}}_{\ast}$ is the local maximizer on $\mathcal{X}$, which satisfies $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}}_{\ast})\leq f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon_{1}\leq c_{\mathrm{gap}}.$ This contradicts condition 1 in Lemma 2.
@@
  • When $c_{\mathrm{gap}}<c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}$, $\varepsilon=c_{\mathrm{gap}}$ holds. If there exists $\bm{x}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ such that $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon=c_{\mathrm{gap}}$, we can choose $\widetilde{\bm{x}}_{\ast}$ such that $\widetilde{\bm{x}}_{\ast}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ since $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}})\geq c_{\mathrm{quad}}\rho_{\mathrm{quad}}^{2}>c_{\mathrm{gap}}$ holds for any $\widetilde{\bm{x}}\in\mathcal{B}_{2}^{b}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$. Furthermore, such $\widetilde{\bm{x}}_{\ast}$ is the local maximizer on $\mathcal{X}$, which satisfies $f(\bm{x}^{\ast})-f(\widetilde{\bm{x}}_{\ast})\leq f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon=c_{\mathrm{gap}}.$ This contradicts condition 1 in Lemma 2.
From the above two arguments, we have $\forall\bm{x}\in\mathcal{X}\setminus\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast}),f(\bm{x}^{\ast})-f(\bm{x})>\varepsilon$. This implies that it is necessary to satisfy $\bm{x}\in\mathcal{B}_{2}(\rho_{\mathrm{quad}};\bm{x}^{\ast})$ under $f(\bm{x}^{\ast})-f(\bm{x})\leq\varepsilon$. ∎
@@
Fix any index set $\mathcal{T}\subset[T]$. Then, when running GP-UCB, we have the following inequality under $\mathcal{A}$:
@@
where the second inequality follows from Theorems 5.3 and 5.4 in [srinivas10gaussian]. ∎
Appendix D Numerical Simulation for Information Gain
In addition to the simple example provided in Figure 1, we empirically confirm the gap between the worst-case (MVR) and GP-UCB’s information gain under the Bayesian assumption. In Figures 2 and 3, we report the average and the quantile of realized information gain with GP-UCB, over 20 different sample paths, generated by changing the random seed, respectively. We conduct experiments under the same settings as in Figure 1 of the main text. We also report the information gain corresponding to the sequence of maximum variance reduction (MVR), following the same setup as Figure 1. In all cases, consistent with Figure 1 in the main text, we observe a noticeable gap in information gain between GP-UCB and MVR.
Figure 2: Average results of information gain under 20 different sample paths. The top row shows the results for the input space $\mathcal{X}=[0,1]$ with lengthscale parameter $\ell=0.1$. The bottom row corresponds to the input space $\mathcal{X}=[0,1]^{2}$ with lengthscale $\ell=0.25$, and from left to right, we use the SE kernel, the Matérn 5/2 kernel, and the Matérn 7/2 kernel. The shaded regions indicate one standard error.
Figure 3: Quantiles of information gain over 20 different sample paths. We report the 10%, 25%, 50%, 75%, and 90% quantiles of the information gain of GP-UCB.